# Lessons Learned from a Unifying Empirical Study of Parameter-Efficient Transfer Learning (PETL) in Visual Recognition

### v1

@@ 2.3 Related work and comparison
The community-wide enthusiasm for PETL has led to multiple survey articles [96, 94, 25]. Meanwhile, several empirical, integrative, and theoretical studies were presented, mostly based on NLP tasks, attempting to provide a holistic understanding. [26, 67] provided unified views to methodologically connect PETL approaches. [8, 15, 28] and [29] empirically compared PETL approaches on NLP and vision tasks, respectively, while [20] offered a theoretical stability and generalization analysis. Accuracy-wise, [8, 15, 28] found that PETL is robust to over-fitting and quite effective in NLP tasks under low-data regimes. This is, however, not the case for vision tasks: [29] showed that representative PETL approaches like LoRA and Adapter cannot consistently outperform either full fine-tuning or linear probing. In terms of why PETL works, [20] framed PETL as sparse fine-tuning and showed that it imposes a regularization by controlling stability; [15, 29] framed PETL as (subspace) optimization; [15] further discussed the theoretical principle inspired by optimal control.
@@
3 Results on Low-Shot Regimes
Dataset. VTAB-1K [99] consists of 19 image classification tasks from three groups. The Natural group comprises natural images captured with standard cameras. The Specialized group contains images captured by specialist equipment for remote sensing and medical purposes. The Structured group evaluates the scene structure comprehension, such as object counting and 3D depth estimation. Following [99], we perform an 80/20 split on the 1000 training images in each task for hyperparameter searching. The reported result (top-1 classification accuracy) is obtained by training on the 1000 training images and evaluating on the original test set. We provide more detail in the Appendix.
Methods. We consider linear probing, full fine-tuning, and 13 PETL methods including 1 prompt-based [39], 5 adapter-based [33, 72, 9, 40, 61], 4 Direct selective [98, 5, 93, 53], and 3 Efficient selective [34, 41]. Please refer to subsection 2.2 for details.
Setup. We employ the ViT-B/16 model [17] pre-trained on ImageNet-21K [13] as the backbone. The prediction head is randomly initialized for each dataset. Images are resized to $224\times 224$. We perform no data augmentation and normalization during training, following [99]. (The exception is CIFAR-100 [46], on which normalization is necessary.) We systematically tune 1) learning rate, 2) weight decay, and 3) approach-specifics like the size of PETL parameters which are often left intact. We set a cap for 3), $\leq 1.5\%$ of ViT-B/16. We also turn the drop path rate on (e.g., 0.1) or off (i.e., 0). A detailed hyper-parameter search grid and additional training details are provided in the Appendix.
Results. As shown in 1(a) and Table 1, PETL methods generally outperform both linear probing and full fine-tuning across datasets. Additionally, under fair hyper-parameter tuning, we surprisingly found that most PETL methods perform similarly as the relative standard deviations (divided by the means) in all three groups are quite low. Simple methods (e.g., Bitfit) and PETL methods originally proposed for NLP (e.g., LoRA and Adapter), which were previously reported as inferior due to un-optimized implementations and hyperparameter tuning, now demonstrate competitive performance with SOTA visual PETL methods. To understand the relative advantages of different approaches, we provide the ranking of PETL methods across different groups in Figure 2, where the element $(i,j)$ in each ranking matrix represents the frequency that method $i$ ranks $j$th in each group. Methods are ordered by their mean ranks (in brackets), and the parameters column indicates the number of trainable parameters in millions. In the natural group, methods with very few trainable parameters perform typically better. Conversely, in the specialized and structured groups, methods with more parameters generally yield better performance. We hypothesize that this performance discrepancy arises from the domain affinity between the pre-trained domain (ImageNet) and the downstream domains. The natural group, sharing a stronger affinity with ImageNet, allows simpler methods like BitFit to adjust the features effectively. In contrast, the specialized and structured groups necessitate more complex methods with more trainable parameters to bridge the domain gap.
Recipes. In low-shot regimes, when the downstream data are similar to the pre-trained data, simple methods (e.g., DiffFit) with decent accuracy and much fewer parameters are preferred. When there is a large domain gap, complex methods with higher accuracy (but more parameters) become competitive. Learning with low-shot data is prone to over-fitting. We find that if the drop path rate — which stochastically drops a transformer block per sample [35] — is set not as default (i.e., nonzero), all the methods can benefit from such a regularization, as shown in Figure 10 in the Appendix.
Table 1: Results on VTAB-1K (19 tasks from 3 groups). We include VPT-Shallow [39] just for reference. Based on the accuracy among PETL, linear probing, and full fine-tuning, we find two task groups (purple and orange).
@@ 3 Results on Low-Shot Regimes
Figure 2: Ranking 15 methods (14 PETL including VPT-Shallow [39], and linear probing) in three groups of VTAB-1K tasks. Element $(i,j)$ is the number of times method $i$ ranks $j$th in each group. Methods are ordered (top-down) by mean ranks (in brackets). The parameters column shows the # of trainable parameters in millions.
@@ 4 PETL Approaches Offer Complementary Information
The previous section shows all PETL methods achieve similar accuracy in various domains. Does it mean all PETL methods learn similar knowledge from the downstream data? As different PETL approaches have different inductive biases — they explicitly select specific parameters to update or insert different modules in various locations — we hypothesize that they may learn differently from the same downstream data.
We start by analyzing their prediction similarity on the same dataset in VTAB-1K. It is expected that their predictions are similar for datasets with very high accuracy, such as Flowers102 (avg 99.1%) and Caltech101 (avg 91.4%). Beyond them, we find that most PETL methods show diverse predictions in other datasets in VTAB-1K. Figure 5 shows the prediction similarities between 14 PETL methods in DTD, Retinopathy, and DMLab, which are from natural, specialized, and structured groups, respectively. In DTD and Retinopathy, most methods show around 20% different predictions while in DMLab, this number goes up to around 35%, even though they achieve similar accuracy.
Such diverse predictions across methods open up the possibility of leveraging their heterogeneity for further improvement. The most straightforward approach is ensemble [23, 48], e.g., average logits over methods. Figure 3 demonstrates the ensemble performance gain over all the PETL methods in each dataset, where we use the worst PETL method as the baseline. Due to the diverse predictions across methods, the ensemble can provide consistent gain.
Additionally, we analyze if PETL methods make similar correct predictions for high-confidence samples and similar mistakes for low-confidence samples. Figure 5 shows the correct prediction overlap for the 5K most confident samples (per method) and the wrong prediction overlap for the 5K least confident samples (per method). For demonstration purposes, we select one method from each PETL category (LoRA, Adapter, SSF) and they are fine-tuned on CIFAR-100 in VTAB-1K. Since they make different predictions in both high and low-confidence regimes, this paves the way for new possibilities of using different PETL methods to generate diverse pseudo-labels for semi-supervised learning [95, 21, 90], domain adaptation [19, 84, 18], continual learning [80, 59, 65] and transfre learning [105, 85, 58].
Figure 3: Ensemble (avg logits) provides consistent gain on most datasets thanks to the diverse prediction.
@@
Figure 4: Prediction similarity analysis: element $(i,j)$ shows the percentage of samples that method $i$ and method $j$ have the same predictions. Although different methods achieve similar overall accuracy, they have quite diverse predictions.
@@
Figure 5: LoRA, Adapter and SSF fine-tuned on CIFAR100 in VTAB-1K. (a): the correct prediction overlap for the 5K most confident samples. (b): the wrong prediction overlap for the 5K least confident samples.
5 Results on Many-Shot Regimes
@@
Recent works in NLP [8] have indicated that PETL methods may not perform as competitively as full fine-tuning when data is abundant. We thus aim to investigate PETL’s performance in many-shot regimes by addressing the following questions: (1) Should we use PETL or full fine-tuning when data is sufficient? (2) How should we adjust the number of trainable parameters for PETL methods in many-shot regimes?
@@
Dataset. We select one representative dataset from each of the natural, specialized, and structured groups in VTAB: (1) CIFAR-100 [46], a natural image dataset comprising 50K training images across 100 classes; (2) RESISC [10], a remote sensing dataset for scene classification with 25.2K training samples across 45 classes; and (3) Clevr-Distance [99, 42], a synthetic image dataset for predicting the depth of the closest object from the camera with 6 depth classes and 70K samples. The reported results are obtained by training on the full training set and evaluating on the original test set.
Setup. The model setup follows the VTAB-1K experiment. We apply horizontal flipping for CIFAR100, horizontal and vertical flipping for Resisc, and no augmentation for Clevr.
@@
Results. In many-shot regimes, with sufficient downstream data, full fine-tuning may catch up and eventually outperform PETL methods. However, from Figure 6, we, found that even in many-shot regimes, PETL can achieve quite comparable results with full fine-tuning, even just using $2\%$ of fine-tuning parameters. (The performance gain, however, quickly diminishes and plateaus after 5% of tunable parameters.) By comparing the results on the domain-close CIFAR-100 and domain-different RESISC and Clevr, we have some further observations. On the one hand, downstream tasks with larger domain gaps suggest the need to update, perhaps many, parameters to obtain high accuracy. With sufficient downstream data, full fine-tuning is less prone to over-fitting and indeed attains a high accuracy. But interestingly, PETL methods, with only $2\sim 5\%$ of tunable parameters, achieve similar accuracy, suggesting that its design principle does offer sufficient effective capacity for the model to learn [100]. On the other hand, downstream tasks with smaller domain gaps suggest that the pre-trained model had learned sufficient knowledge about them; fully fine-tuning it thus risks washing such knowledge away. In fact, we found that PETL notably outperforms full fine-tuning on CIFAR-100, suggesting it as a more robust transfer learning algorithm for downstream tasks.
@@
Figure 6: PETL accuracy in many-shot regimes, with different parameter sizes on three datasets from different domains. Even a few trainable parameters allow the model to have sufficient capacity to learn from full data.
6 Why Do PETL Methods Work?
Within 19 VTAB-1K tasks, we see two cases: (1) Full fine-tuning outperforms linear probing. Since linear probing reflects the pre-trained feature quality for downstream tasks, case (1) suggests the necessity to update the backbone to address the gap between pre-trained and downstream domains. (2) Linear probing surpasses full fine-tuning, suggesting the pre-trained features are good enough (at least in a low-shot scenario). Recklessly updating them with limited data may risk over-fitting.
Figure 7 (a-b) summarizes the low-shot accuracy comparison based on the categorization above. Specifically, each line corresponds to one task. Linear probing, PETL, and fine-tuning are located in order, from left to right, to reflect their tunable parameter sizes. PETL’s superior performance in both cases showcases its capacity to learn and its regularization role to prevent over-fitting.
@@ 6 Why Do PETL Methods Work?
Figure 7: (a): VTAB-1K tasks in case 1, PETL $>$ full $>$ linear . (b) VTAB-1K tasks in case 2, PETL $>$ linear $>$ full. (c) RESISC & Clevr in case 1 with sufficient data, PETL $\approx$ full $>$ linear. (d) CIFAR in case 2 with sufficient data, PETL $>$ full $>$ linear. Within each figure, left for linear, middle for PETL, and right for full.
7 Empirical Results on Robustness to Distribution Shifts
@@
Large pre-trained models such as CLIP [73] and ALIGN [38] have demonstrated unprecedented accuracy across a range of data distributions when performing zero-shot inference. However, recent studies [92, 73] have shown that fine-tuning on downstream data, while significantly boosting performance on the target distribution, often compromises the model’s robustness to distribution shifts. Given that PETL only updates a limited number of parameters in the model, we investigate whether PETL can offer a more robust alternative to full fine-tuning for pre-trained models.
@@
Dataset. We use 100-shot ImageNet-1K as our target distribution, with each class containing 100 images. Following [92], we consider 4 natural distribution shifts from ImageNet: ImageNet-V2 [76], a new ImageNet test set collected with the original labeling protocol; ImageNet-R [31], renditions for 200 ImageNet classes; ImageNet-S [22], sketch images for 1K ImageNet classes; ImageNet-A [32], a test set of natural images misclassified by a ResNet-50 [27] for 200 ImageNet classes.
@@
Setup. We focus on the CLIP ViT-B/16 model, which comprises a visual encoder and a text encoder, pre-trained via contrastive learning on image-text pairs. Following [92], we add an FC layer as the prediction head with zero-initialized bias and initialize weights using the class label text embedded by the text encoder. Subsequently, we discard the text encoder and apply PETL methods to the visual encoder, fine-tuning only the PETL modules and the head.
@@
Results. As shown in Table 2, while some PETL methods may not surpass full fine-tuning on the target distribution, they consistently demonstrate more robust performance on OOD data. This is likely because PETL updates only a small fraction of the parameters, thus preserving the robust features of the foundation models. Given the similar target distribution performance, should we blindly use PETL methods for more robustness?
@@
Weight-space ensembles (WiSE) for PETL. WiSE [92], which linearly interpolates the full fine-tuned and original models, is a popular fine-tuning approach to enhance robustness. We explore whether WiSE can enhance the robustness of PETL. To apply WiSE to PETL, we first linearly interpolate the prediction head with a mixing coefficient $\alpha$. For direct selective tuning methods (e.g.BitFit), we directly interpolate with the original model. Since most Adapter-based methods have residual connections, we can multiply the adapter modules with $\alpha$ to control their strengths. A similar approach can be applied to efficient selective methods (e.g.LoRA) as they learn additive residuals to the original parameters. As shown in 1(b), WiSE improves both the robustness and the in-distribution performance of PETL methods. Interestingly, even though full fine-tuning is generally less robust than PETL methods, applying WiSE allows it to achieve better performance in both in-distribution and OOD data, which suggests a promising research direction for robust PETL.
@@
100-shot ImageNet 74.95 75.27 74.80 75.01 75.62 76.50 76.26 76.56 74.69
Avg. OOD Acc 42.50 55.44 55.91 56.12 56.12 56.21 54.65 55.93 56.12
Table 2: The Avg. OOD Acc shows the average performance for ImageNet-(V2, S, R, A).

### v2

@@ 2.3 Related work and comparison
The community-wide enthusiasm for PETL has led to multiple survey articles (Yu et al., 2023a; Xin et al., 2024; Han et al., 2024). Meanwhile, several empirical, integrative, and theoretical studies were presented, mostly based on NLP tasks, attempting to provide a holistic understanding. (He et al., 2021a; Mao et al., 2021) provided unified views to methodologically connect PETL approaches. (Chen et al., 2022a; Ding et al., 2023; He et al., 2021b) and (He et al., 2022a) empirically compared PETL approaches on NLP and vision tasks, respectively, while (Fu et al., 2023) offered a theoretical stability and generalization analysis. Accuracy-wise, (Chen et al., 2022a; Ding et al., 2023; He et al., 2021b) found that PETL is robust to over-fitting and quite effective in NLP tasks under low-data regimes. This is, however, not the case for vision tasks: (He et al., 2022a) showed that representative PETL approaches like LoRA and Adapter cannot consistently outperform either full fine-tuning or linear probing. In terms of why PETL works, (Fu et al., 2023) framed PETL as sparse fine-tuning and showed that it imposes a regularization by controlling stability; (Ding et al., 2023; He et al., 2022a) framed PETL as (subspace) optimization; (Ding et al., 2023) further discussed the theoretical principle inspired by optimal control.
@@
3 How Do PETL Methods Perform in Low-Shots Regime?
@@
Table 1: Results on VTAB-1K (19 tasks from 3 groups). Based on the accuracy among PETL, linear probing, and full fine-tuning, we find two task groups (purple and orange), as discussed in section 6.
@@
Figure 2: Ranking frequency of 15 methods (14 PETL + linear probing) for three groups in VTAB-1K. Element $(i,j)$ is the number of times method $i$ ranks $j^{th}$ in each group. Methods are ordered by mean ranks (in brackets). The parameters column shows the # of trainable parameters in millions. More details are in Appendix C.
Dataset. VTAB-1K (Zhai et al., 2019) consists of 19 image classification tasks from three groups. The Natural group comprises natural images captured with standard cameras. The Specialized group contains images captured by specialist equipment for remote sensing and medical purposes. The Structured group evaluates the scene structure comprehension, such as object counting and 3D depth estimation. Following Zhai et al. (2019), we perform an 80/20 split on the 1000 training images in each task for hyperparameter tuning. The reported top1 accuracy is obtained by training on the 1000 training images and evaluating on the original test set.
Methods. We consider linear probing, full fine-tuning, and 14 PETL methods including 2 prompt-based (Jia et al., 2022), 5 adapter-based (Houlsby et al., 2019; Pfeiffer et al., 2021; Chen et al., 2022b; Jie & Deng, 2022; Luo et al., 2023), 4 Direct selective (Zaken et al., 2022; Basu et al., 2023; Xie et al., 2023; Lian et al., 2022), and 3 Efficient selective (Hu et al., 2021; Jie & Deng, 2023). Please refer to subsection 2.2 for details.
Setup. We employ the ViT-B/16 model (Dosovitskiy et al., 2020) pre-trained on ImageNet-21K (Deng et al., 2009) as the backbone. The prediction head is randomly initialized for each dataset. Images are resized to $224\times 224$. We systematically tune 1) learning rate, 2) weight decay, and 3) approach-specifics like the size of PETL parameters which are often left intact. We set a cap for 3), $\leq 1.5\%$ of ViT-B/16. We also turn the drop path rate (Huang et al., 2016) on (e.g., 0.1) or off (i.e., 0). A detailed hyperparameter search grid and additional training details are provided in Appendix A.1.
Results. As shown in 1(a) and Table 1, PETL methods generally outperform both linear probing and full fine-tuning across datasets. Additionally, under fair hyper-parameter tuning, we surprisingly found that most PETL methods perform similarly as the relative standard deviations (divided by the means) in all three groups are quite low. Simple methods (e.g., Bitfit) and PETL methods originally proposed for NLP (e.g., LoRA and Adapter), which were previously reported as inferior due to un-optimized implementations and hyperparameter tuning, now demonstrate competitive performance with SOTA visual PETL methods. To understand the relative advantages of different approaches, we provide the ranking frequency of PETL methods across different groups in Figure 2, where the element $(i,j)$ in each ranking matrix represents the frequency that method $i$ ranks $j^{th}$ in each group. Methods are ordered by their mean ranks (in brackets), and the parameters column indicates the number of trainable parameters in millions. In the natural group, simpler methods with fewer trainable parameters—such as DiffFit and Fact-TT—offer a cost-effective solution without compromising performance. Conversely, in the specialized and structured groups, methods with more parameters generally yield better performance. We hypothesize that this performance discrepancy arises from the domain affinity between the pre-trained domain (ImageNet) and the downstream domains. The natural group, sharing a stronger affinity with ImageNet, allows simpler methods like BitFit to adjust the features effectively. In contrast, the specialized and structured groups necessitate more complex methods with more trainable parameters to bridge the domain gap.
Recipes. In low-shot regimes, when the downstream data are similar to the pre-trained data, simple methods (e.g., DiffFit) with decent accuracy and much fewer parameters are preferred—aligning with Occam’s razor. When there is a substantial domain gap, complex methods with higher accuracy (but more parameters) become competitive. Learning with low-shot data is prone to over-fitting. We find that if the drop path rate — which stochastically drops a transformer block per sample (Huang et al., 2016) — is set not as default (i.e., nonzero), all the methods can benefit from such a regularization, as shown in Figure 10 in the Appendix.
@@ 4 PETL Approaches Offer Complementary Information
Figure 3: Prediction similarity analysis: element $(i,j)$ shows the percentage of samples that method $i$ and method $j$ have the same predictions. Although different methods achieve similar accuracy, they have diverse predictions. Details in Appendix C.
@@
Figure 4: LoRA, Adapter and SSF fine-tuned on CIFAR100(VTAB-1K). (a): correct prediction overlap for the 5K most confident data. (b): wrong prediction overlap for the 5K least confident data. Details in Appendix C.
The previous section demonstrates that all PETL methods perform similarly across various domains. Given that different PETL methods are trained on the same downstream data using the same backbone and achieve comparable accuracy, one might expect them to learn similar knowledge from the data, resulting in similar predictions. Contrary to this expectation, our findings below reveal that different PETL methods acquire distinct and complementary knowledge from the same downstream data, even when built upon the same backbone, leading to diverse predictions.
Figure 5: Ensemble (avg logits) provides consistent gain on most datasets thanks to the diverse prediction. Details in Appendix C.
We start by analyzing their prediction similarity on the same dataset in VTAB-1K. It is expected that their predictions are similar for datasets with very high accuracy, such as Flowers102 (avg 99.1%) and Caltech101 (avg 91.4%). Beyond them, we find that most PETL methods show diverse predictions in other datasets in VTAB-1K. Figure 4 shows the prediction similarities between 14 PETL methods in DTD, Retinopathy, and DMLab, which belong to natural, specialized, and structured groups, respectively. In DTD and Retinopathy, most methods differ in about 20% of their predictions, while in DMLab, this difference increases to approximately 35%, even though they achieve similar accuracies. This prediction diversity may be attributed to the different inductive biases (Neyshabur et al., 2014) of PETL methods — they explicitly select specific parameters to update or insert different modules at various locations within the model. More analyses and details about Figure 4 are offered in Appendix C.
Such diverse predictions across methods open up the possibility of leveraging their heterogeneity for further improvement. The most straightforward approach is ensemble (Gontijo-Lopes et al., 2021), e.g., average logits over methods. Figure 5 demonstrates the ensemble performance gain over all the PETL methods in each dataset, where we use the worst PETL method as the baseline. Due to the diverse predictions across methods, the ensemble can provide consistent gain.
Also, we analyze if PETL methods make similar correct predictions for high-confidence samples and similar mistakes for low-confidence samples. Figure 4 shows the correct prediction overlap for the 5K most confident samples (per method) and the wrong prediction overlap for the 5K least confident samples (per method). For demonstration purposes, we select one method from each PETL category (LoRA, Adapter, SSF) and they are fine-tuned on CIFAR-100 in VTAB-1K. Methods within the same category also show diverse predictions (Appendix C). Since they make different predictions in both high and low-confidence regimes, this paves the way for new possibilities of using different PETL methods to generate diverse pseudo-labels for semi-supervised learning Yang et al. (2022); Gan & Wei ; Wang et al. (2022b), domain adaptation Farahani et al. (2021); Tang et al. (2024); Englert et al. (2024), continual learning Shim et al. (2021); Lomonaco et al. (2022); Mai et al. (2021) and transfre learning Zhuang et al. (2020); Tu et al. (2024); Liu et al. (2022c).
5 How Do PETL Methods Perform in Many-Shot Regime?
@@
Recent works in NLP Chen et al. (2022a) have indicated that PETL methods may not perform as competitively as full fine-tuning when data is abundant. We thus aim to investigate PETL’s performance in many-shot regimes by addressing the following questions: (1) Should we use PETL or full fine-tuning when data is sufficient? (2) How should we adjust the number of trainable parameters for PETL methods in many-shot regimes?
@@
Dataset. We select one representative dataset from each of the natural, specialized, and structured groups in VTAB: (1) CIFAR-100 Krizhevsky et al. (2009), a natural image dataset comprising 50K training images across 100 classes; (2) RESISC Cheng et al. (2017), a remote sensing dataset for scene classification with 25.2K training samples across 45 classes; and (3) Clevr-Distance Zhai et al. (2019); Johnson et al. (2017), a synthetic image dataset for predicting the depth of the closest object from the camera with 6 depth classes and 70K samples. The reported results are obtained by training on the full training set and evaluating on the original test set.
Setup. The model setup follows the VTAB-1K experiment. More details about setup and hyperparameter search are provided in Appendix A.
@@
Results. In many-shot regimes, with sufficient downstream data, full fine-tuning may catch up and eventually outperform PETL methods. However, from Figure 6, we found that even in many-shot regimes, PETL can achieve comparable results with full fine-tuning, even just using $2\%$ of fine-tuning parameters. (The performance gain, however, quickly diminishes and plateaus after 5% of tunable parameters.) By comparing the results on the domain-close CIFAR-100 and domain-different RESISC and Clevr, we have some further observations. On the one hand, downstream tasks with larger domain gaps suggest the need to update, perhaps many, parameters to obtain high accuracy. With sufficient downstream data, full fine-tuning is less prone to over-fitting and indeed attains a high accuracy. But interestingly, PETL methods, with only $2\sim 5\%$ of tunable parameters, achieve similar accuracy, suggesting that its design principle does offer sufficient effective capacity for the model to learn Zhang et al. (2021). On the other hand, downstream tasks with smaller domain gaps suggest that the pre-trained model had learned sufficient knowledge about them; fully fine-tuning it thus risks washing such knowledge away. In fact, we found that PETL notably outperforms full fine-tuning on CIFAR-100, suggesting it as a more robust transfer learning algorithm for downstream tasks.
@@
Figure 6: PETL accuracy in many-shot regimes, with different parameter sizes (X-axis) on three datasets from different domains. Even 2%-5% trainable parameters allow the models to have sufficient capacity to learn from full data. (Details are in Appendix C)
6 Why Do PETL Methods Work?  [^2] Our intention is not to offer a definitive conclusion about why PETL works. As discussed in subsection 2.3, there is currently no universally agreed-upon explanation for the effectiveness of PETL. We hope our empirical findings will contribute to the ongoing efforts to understand the underlying principles of PETL methods.
Putting together section 3 and section 5, we identify two distinct patterns regarding the performance among linear probing, full fine-tuning, and PETL. Within 19 VTAB-1K tasks, we see: (1) Full fine-tuning outperforms linear probing. As linear probing reflects the pre-trained feature quality for downstream tasks, case (1) suggests the necessity to update the backbone to close the gap between pre-trained and downstream domains. (2) Linear probing surpasses full fine-tuning, suggesting the pre-trained features are good enough (at least in a low-shot scenario). Recklessly updating them may risk over-fitting. Figure 7 (a-b) summarizes the low-shot accuracy comparison based on the categorization above; each line corresponds to one task. Linear probing, PETL, and fine-tuning are located in order, from left to right, to reflect their tunable parameter sizes. PETL’s superiority in both cases showcases its capacity to learn and its regularization role to prevent over-fitting.
@@
Figure 7: (a): VTAB-1K tasks in case (1) , PETL $>$ full $>$ linear . (b) VTAB-1K tasks in case (2), PETL $>$ linear $>$ full. (c) RESISC & Clevr in case (1) with enough data, PETL $\approx$ full $>$ linear. (d) CIFAR in case (2) with enough data, PETL $>$ full $>$ linear. Within each figure, left for linear, middle for PETL, and right for full. More details are in Appendix C.
7 Are PETL methods more robust to distribution shifts?
@@
Large pre-trained models such as CLIP Radford et al. (2021) and ALIGN Jia et al. (2021) have demonstrated unprecedented accuracy across a range of data distributions when performing zero-shot inference. However, recent studies Wortsman et al. (2022); Radford et al. (2021) have shown that fine-tuning on downstream data, while significantly boosting performance on the target distribution, often compromises the model’s robustness to distribution shifts. Given that PETL only updates a limited number of parameters in the model, we investigate whether PETL can offer a more robust alternative to full fine-tuning for pre-trained models.
@@
Dataset. We use 100-shot ImageNet-1K as our target distribution, with each class containing 100 images. Following Wortsman et al. (2022), we consider 4 natural distribution shifts from ImageNet: ImageNet-V2 Recht et al. (2019), a new ImageNet test set collected with the original labeling protocol; ImageNet-R Hendrycks et al. (2021a), renditions for 200 ImageNet classes; ImageNet-S Gao et al. (2023), sketch images for 1K ImageNet classes; ImageNet-A Hendrycks et al. (2021b), a test set of natural images misclassified by a ResNet-50 He et al. (2015) for 200 ImageNet classes.
@@
Setup. We focus on the CLIP ViT-B/16 model, which comprises a visual encoder and a text encoder, pre-trained via contrastive learning on image-text pairs. Following Wortsman et al. (2022), we add an FC layer as the prediction head with zero-initialized bias and initialize weights using the class label text embedded by the text encoder. Subsequently, we discard the text encoder and apply PETL methods to the visual encoder, fine-tuning only the PETL modules and the head. More details about the CLIP model and experiment setup can be found in Appendix A.1.
@@
Results. As shown in Table 2, while some PETL methods may not surpass full fine-tuning on the target distribution, they consistently demonstrate more robust performance on distribution shift data. This is likely because PETL updates only a small fraction of the parameters, thus preserving the robust features of the foundation models. Given the similar target distribution performance, should we blindly use PETL methods for more robustness?
@@
Weight-space ensembles (WiSE) for PETL. WiSE (Wortsman et al., 2022), which linearly interpolates the full fine-tuned and original models, is a popular fine-tuning approach to enhance robustness. We explore whether WiSE can enhance the robustness of PETL. To apply WiSE to PETL, we first linearly interpolate the prediction head with a mixing coefficient $\alpha$. For direct selective tuning methods (e.g.BitFit), we directly interpolate with the original model. Since most Adapter-based methods have residual connections, we can multiply the adapter modules with $\alpha$ to control their strengths. A similar approach can be applied to efficient selective methods (e.g.LoRA) as they learn additive residuals to the original parameters. As shown in 1(b) (more results in Appendix C), WiSE improves both fine-tuning and distribution shift performance of PETL methods. Interestingly, even though full fine-tuning is generally less robust than PETL methods, applying WiSE allows it to achieve better performance in both target distribution and distribution shift data, which suggests a promising research direction for robust PETL.
@@
100-shot ImageNet 75.0 75.27 74.8 75.0 75.6 76.5 76.3 76.6 74.7
Avg. distribution shift Acc 42.5 55.4 $\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}(12.9){\uparrow}$ 55.9 $\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}(13.4){\uparrow}$ 56.9 $\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}(14.4){\uparrow}$ 56.1 $\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}(13.6){\uparrow}$ 56.2 $\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}(13.7){\uparrow}$ 54.7 $\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}(12.2){\uparrow}$ 55.9 $\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}(13.4){\uparrow}$ 56.1 $\color[rgb]{0,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,0}(13.6){\uparrow}$
[…]